The ACMAL Logbook system will be down on the afternoon of Sunday 12/13. After hours users, who will be using the instruments during this time, please email me. I will need to know your time and the index you will be using.
The sputter coater needs a new target. I’m working with our supplier and should have a replacement soon. Until then we can coat your samples with carbon. Please contact Aleister Kerr or myself if you need any samples coated.
The FIB repairs are complete. The new tip has been installed and it is ready for use.
The FIB tip will be arriving next week. The Hitachi technician will be installing it shortly thereafter. I’ll send out an update when the repair is complete.
This is a reminder that ACMAL facilities will be CLOSED on Thursday and Friday, November 26th & 27th. The building and ACMAL labs will be open Monday, Tuesday & Wednesday for restricted users (with no keys or tap access). Unrestricted users who have building and lab tap access are free to use the equipment anytime. If you are a new user please email us to arrange for supervision in case you have a problem.
ACMAL staff is available by appointment. Please reach out by email or phone if you have any questions or concerns.
The ACMAL labs will continue to be open for faculty, staff, graduate students, and undergraduates who’s research cannot be completed from home. If you require access, and have not already let me or Valentina know, please complete the ACMAL Building Access Form.
Prior to using the facility, review the ACMAL COVID-19 Standard Operating Procedure. Our staff will continue to sanitize the labs daily. However, we are now asking all users to complete a touchpoint cleaning after they use the lab.
Please let me know if you have any questions,
The FIB is down until we can get the tip replaced. I am working with Hitachi on a timeline.
The EDS system on the Field Emission is now back up and running.
The EDS system on the Field Emission is down. The power supply on the computer that interfaces with the EDS has failed. We have ordered a new power supply and will get it installed as soon as possible.
The EDS system on the Field Emission is back up and running.